OpenStudy (lgbasallote):

how to solve for \(T_m\) using a calculator? \[\huge \frac{\ln (\frac{40 - T_m}{70-T_m})}{5} = \frac{\ln (\frac{0.5}{70-T_m})}{35}\]

OpenStudy (anonymous):

Is it a simple scientific calculator?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

isolate Tm (on the right hand side equation)

OpenStudy (lgbasallote):

hmm?

OpenStudy (anonymous):

u will have a equation like: Tm=f(Tm) enter an initial guess to save it in the Ans then enter f(Tm) in ur calculator with Ans instead of Tm then push the Ans Button repeatedly to converge

OpenStudy (lgbasallote):

wahhh i dont get you...go slow please :(

OpenStudy (phi):

you could use newton's method \[ y_{n+1}= y_n - \frac{f(y_n)}{f'(y_n)} \]

OpenStudy (lgbasallote):

wahh? o.O

OpenStudy (lgbasallote):

lol i only need to know how to input this in calculator :C

OpenStudy (phi):

I believe you can only find a solution using a numerical technique, an iterative search that converges on an answer. A bit painful using a calculator. If you just want the solution, use wolfram.

OpenStudy (lgbasallote):

but i need to learn howto use my calculator...hmm will you just tell me how to isolate tm then?

OpenStudy (phi):

you can get to a form like \[ 2(40-T_m)^7=(7-T_m)^6 \] but you can not isolate Tm...

OpenStudy (lgbasallote):

hmm...where did 2 come from again?

OpenStudy (phi):

*70 the 2 from 0.5

OpenStudy (lgbasallote):

oh og course

OpenStudy (lgbasallote):

of course*

OpenStudy (lgbasallote):

@phi are you familiar with the shift + solve in scientific calculators?

OpenStudy (phi):

No, but I assume it can solve equations numerically?

OpenStudy (lgbasallote):

i think so..if i can figure it out...

OpenStudy (phi):

maybe this will help for convenience, let y= 40-x, so the problem can be written as \[2y^7= (y+30)^6 \] \[ 2^{(1/6)}y^{(7/6)}-y-30 = 0 \] now you find the roots of this expression, and then find x= 40-y

OpenStudy (lgbasallote):

hmm i did not get that lol sorry...i'll try doing trial and error on my calculator now

OpenStudy (phi):

what calculator are you using? so I can look at its user manual

OpenStudy (lgbasallote):

canon f-788dx